OpenStudy (briggles1999):

write an equation in point slope form for the line through the given point with the given slope. (8,3) m=6

OpenStudy (briggles1999):

y+3=6(x-8) y-3=6(x-8) y-3=6(x+8) y+3=6(x+8)